OpenStudy (anonymous):

What is the average of all of the integers from 13 to 37?

OpenStudy (apoorvk):

sum them all up with the AP sum formula... then divide by the no. of nos. to get the mea, or the 'avg.'

OpenStudy (experimentx):

(13+37)/2 .. if it is symmetric on some no

OpenStudy (experimentx):

I thought @Tributized might be right .. if you can constrct (24+25+26)/3 = 25

OpenStudy (anonymous):

i deleted it because u were teaching the asker ._. i don't wanna give the answer yet lol

OpenStudy (experimentx):

(23+24+25+26+27)/5 = 25 (13+ ... +23+24+25+26+27+ ,,,,, 37)/(37-13) = 25

OpenStudy (experimentx):

no .. i just came up with a hunch .. i haven't calculated yet ... this is just interesting ... to predict result without calculating

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

oh ok. so the answer is 25. :D

OpenStudy (experimentx):

seriously, i haven't checked it yet .... just a prediction

OpenStudy (experimentx):

everybody is giving me medals .. guess i was right or what :D

OpenStudy (turingtest):

\[\frac12[\sum_{i=1}^{37}i-\sum_{i=1}^{12}i]\]eh? may not be the most efficient way

OpenStudy (anonymous):

yep ur right. since im right :D

OpenStudy (anonymous):

The average (or mean) is the same as the median with a series of numbers which differentiate from each other by the same number (in this case, 1)

OpenStudy (anonymous):

(13+37)/2 = (14+36)/2 = (15+35)/2 = ... = (24+26)/2 = 25/1 = 25, you only needed to calculate (13+37)/2 =25. The same is true for any sum of consecutive integers.

OpenStudy (anonymous):

Turing, why dividing by 2?

OpenStudy (turingtest):

oops divide by 37-12

OpenStudy (turingtest):

right?

OpenStudy (anonymous):

the sum is \[\S_{n}=\frac{n}{2}(a_{1}+a_{n})\] the average would be \[\frac{\S_{n}}{a_{n}-a_{1}}=\frac{\frac{a_{n}-a_{1}}{2}(a_{1}+a_{n})}{a_{n}-a_{1}}=\frac{\frac{37-13}{2}(37+13)}{37-13}=\frac{37+13}{2}\]

OpenStudy (anonymous):

13 through 37 is the same as 12(25) = 300 ( there are 25 numbers to add ) add to this the numbers 1 through 25 The formula for this is: (1/2) x 25 x 26 = 325 300 + 325 = 625 625/25 = 25

OpenStudy (anonymous):

I did this \[ (37 \times 38/2 - 12 \times 13/2)/(37-12) = 25\]

OpenStudy (anonymous):

yeah that's easier lol

OpenStudy (turingtest):

oh jeez, I added in my calculator at least that earlier formula is right then with the modification mentioned by FFM\[\frac1{25}[\sum_{i=1}^{37}i-\sum_{i=1}^{12}i]\]